Support department cost allocation—reciprocal services method

Davis Snowflake & Co. produces Christmas stockings in its Cutting and Sewing departments. The Maintenance and Security departments support the production of the stockings. Costs from the Maintenance Department are allocated based on machine hours, and costs from the Security Department are allocated based on asset value. Information about each department is provided in the following table:

Maintenance
Department
Security
Department
Cutting
Department
Sewing
Department
Machine hours 800    2,000 7,200    10,800
Asset value $2,000    $1,670 $2,500 $5,500
Department cost $36,000    $16,000 $64,000 $82,000

Determine the total cost of each production department after allocating all support department costs to the production departments using the reciprocal services method.

Cutting
Department
Sewing
Department
Production departmentsʼ total costs $ $
Answer #1

Solution

Davis Snowflake & Co

Determination of the total cost of each production department after allocation of all support department costs to the production departments using the reciprocal services method:

Maintenance

Security

Cutting

Sewing

Department costs

$36,000

$16,000

$64,000

$82,000

Allocation of Security cost

$4,000

$5,000

$11,000

Allocation of Maintenance Cost

($40,000)

$4,000

$14,400

$21,600

($20,000)

Total Costs

0

0

$83,400

$114,600

Computations:

Allocation ratios of service departments –

Security Department to other departments –

Total asset value = 2,000 + 2,500 + 5,500 = 10,000

To maintenance = 2,000/10,000 = 20%

To Cutting = 2,500/10,000 = 25%

To Sewing = 5,500/10,000 = 55%

Maintenance to other departments –

Total machine hours = 2,000 + 7,200 + 10,800 = 20,000

To security = 2,000/20,000 = 10%

To Cutting = 7,200/20,000 = 36%

To Sewing = 10,800/20,000 = 54%

Assume total Maintenance department cost = M

Assume total Security department cost = S

Maintenance department cost allocation –

Total maintenance cost be = 36,000 + 0.2 S

S = 16,000 + 0.1M

Substituting the values of S for total maintenance cost,

M = 36,000 + 0.2 (16,000 + 0.1M)

M = 36,000 + 3,200 + 0.02M

0.98M = 39,200

M = $40,000

Now, substituting value of M = 40,000 for total security cost,

S = 16,000 + 0.1 x 40,000 = $20,000

Maintenance department cost allocation to Security Department = 10% x 40,000 = $4,000

Maintenance department cost allocation to Cutting Department = 36% x 40,000 = $14,400

Maintenance department cost allocation to Sewing Department = 54% x 40,000 = $21,600

Maintenance Department to other departments

Security

Cutting

Sewing

proportion

10%

36%

54%

total cost $40,000

$4,000

$14,400

$21,600

Security department cost allocation to Maintenance Department = 20% x 20,000 = $4,000

Security to Cutting = 25% x 20,000 = $5,000
Security to Sewing = 55% x 20,000 = $11,000

Security Department to other departments

Maintenance

Cutting

Sewing

proportion

20%

25%

55%

total cost $20,000

$4,000

$5,000

$11,000